OpenStudy (anonymous):

Can someone help me find the derivative of the f(x): (8)/(sqrt x-2), using the difference quotient

OpenStudy (anonymous):

Let f(x)=8, and g(x)=(sqrt(x-2) \[(f'g-g'f)/g^2\] f(x)=8. f'(x)=0 \[g(x)=\sqrt{x-2}\] \[d/dx (f(x)/g(x)) = \frac{g'(x)f}{g^2}\]

OpenStudy (anonymous):

can someone help me using the difference quoteint: (f(x+h)-f(x))/h

OpenStudy (anonymous):

when you say "difference quotient" you mean the definition yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[lim_{h->0}\frac{f(x+h)-f(x)}{h}\]

OpenStudy (anonymous):

\[f(x)=\frac{8}{\sqrt{x-2}}\] yes?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok for ease of writing i am going to skip writing the limit each time. we will take the limit at the end.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and also just remember that we are going to have to divide by h at some point. are main job is to simplify \[f(x+h)-f(x)=\frac{8}{\sqrt{x+h-2}}-\frac{8}{\sqrt{x-2}}\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so we just subtract and get \[\frac{8\sqrt{x-2}-8\sqrt{x+h-2}}{\sqrt{x+h-2}\times\sqrt{x-2}}\]

OpenStudy (anonymous):

now we can pull the 8 right in front of the limit, but we could have done that at the beginning. the gimmick now is to multiply top and bottom by the conjugate of the numerator, namely \[\sqrt{x-2}+\sqrt{x-2+h}\]

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

when you do that , your numerator will be \[x-2-(x+h-2)=-h\]

OpenStudy (anonymous):

which is nice because this is the h that will cancel with the h in the denominator

OpenStudy (anonymous):

you denominator will be just what you get, and ugly mess namely \[h\times\sqrt{x+h-2}\times \sqrt{x-2}(\sqrt{x-2}+\sqrt{x+h-2})\]

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

but now you can cancel the h from the numerator and denominator. you numerator will just be -1 (or -8 when you bring the 8 back) and the denominator will be what i wrote about without the in it namely \[\sqrt{x+h-2}\times \sqrt{x-2}(\sqrt{x-2}+\sqrt{x+h-2})\]

OpenStudy (anonymous):

how do we clean up the denominator

OpenStudy (anonymous):

well now you can take the limit as h->0 because you will not get 0/0 if you do. replace h by zero now

OpenStudy (anonymous):

in other words take \[lim_{x->0}\frac{-8}{\sqrt{x+h-2}\times \sqrt{x-2}(\sqrt{x-2}+\sqrt{x+h-2}})\]

OpenStudy (anonymous):

no more zero over zero. we solved that problem when we canceled the h in the denominator with the h in the numerator. so now wherever you see an "h" replace it by 0

OpenStudy (anonymous):

The 8 that we took out now gets placed back in the numerator

OpenStudy (anonymous):

actually look smart and put it right outside the limit sign

OpenStudy (anonymous):

that really should be step 1

OpenStudy (anonymous):

so what would be the derivative, if you dont mind

OpenStudy (anonymous):

but we do put the 8 back into the num inorder to get the derivative right?

OpenStudy (anonymous):

the first line should look like \[8 lim_{h->0}\frac{1}{h} \times (\frac{1}{\sqrt{x+h-2}}-\frac{1}{\sqrt{x-2}})\]

OpenStudy (anonymous):

take the 8 right out front. yes we put it at the end

OpenStudy (anonymous):

so now your last line should look like \[8\times \frac{-1}{\sqrt{x-2}\times \sqrt{x-2}(\sqrt{x-2}+\sqrt{x-2})}\]

OpenStudy (anonymous):

i just replaced all h's in the denominator by 0

OpenStudy (anonymous):

in the expression about 11 lines up

OpenStudy (anonymous):

I got it thank you

OpenStudy (anonymous):

now some algebra to get \[\frac{-8}{(x-2)(2\sqrt{x-2})}=\frac{-4}{(x-2)\sqrt{x-2}}\]

OpenStudy (anonymous):

I might be asking to much of you, but could you help me understand how to use the alternative formula for derivative, namely : f prime (x)= f(z)-f(x)/x-x

OpenStudy (anonymous):

z-x on the denominator

OpenStudy (anonymous):

perhaps of x^2-3x+4

OpenStudy (anonymous):

you mean \[lim_{z->x}\frac{f(z)-f(x)}{z-x}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

this is really just an exercise in algebra for your example. it should go quick. especially since we know the answer is \[f'(x)=2x-3\]

OpenStudy (anonymous):

trick of course will be to factor an (z-x) out of the numerator and cancel.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

we get \[\frac{z^2-3z+4-(x^2-3x-4)}{z-x}\] \[=\frac{z^2-x^2-3z+3x}{z-x}\]

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

now look to factor. first term is diffeernce of two squares, second to obviously have common factor of 3

OpenStudy (anonymous):

\[\frac{(z-x)(z+x)-3(z-x)}{z-x}\]

OpenStudy (anonymous):

or \[\frac{(z-x)(z+x-3)}{z-x}\]

OpenStudy (anonymous):

cancel the z - x and get \[z+x-3\] and now take the limit as z ->x by simply replacing the x by a z to get \[2x-3\] and we knew

OpenStudy (anonymous):

i meant of course "by simply replacing the z by an x"

OpenStudy (anonymous):

in every one of these problems, you goal is to factor and cancel

OpenStudy (anonymous):

okay, i was confused about what to do with that form

OpenStudy (anonymous):

either factor out an "h" in the first form, or factor out a (z-x) in the second

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

they are identical with h replacing z - x

OpenStudy (anonymous):

yw